mass cyclist together with bike 90kg calculate increase in k.E speed increase 6km/h to 12km/h

Grade:

3 Answers

Sharandeep singh
11 years ago

initial kinetic energy is 1/2*m*v^2

1/2 * 90 * 36 = 1620

final k.e. 1/2 * 90 * 144 =6480

increse = 6480 - 1620 = 4860 j

simmie jaglan
11 years ago

given mass=90

u=(6*5)/18m/s   v=(12*5)/18m/s

chamge in ke=1/2*mass*(v*v-u*u)

=375J

Tanmay Warwadkar
11 years ago

 Initial K.E.= 1/2 mv2 = 1/2×90×(6)2 = 1620

 Final K.E.= 1/2 mv2 = 1/2×90×(12)2 = 6480

Increase in K.E.= Final K.E.- Initial K.E.= 6480-1620=4860.